OpenStudy (kmullis6):

Find the mean, median and mode for the following data: 11, 25, 25, 26, 27, 28 Choose the correct order for listing these measures of central tendency from least to greatest. a) mean, mode, median b) median, mode, mean c) mean, median, mode d) mode, median, mean

OpenStudy (anonymous):

a)

OpenStudy (anonymous):

Mean: (11+25+25+26+27+28)/6 = 28.4 Mode: Most occurring = 25 Median: it's already in ascending order, find middle term = (25+26)/2 = 25.5 B